A stationary bomb explodes into two parts of masses 3kg and 1kg. The total kinetic energy of the two parts after explosion is 2400 J. Then the K.E. of the smaller part is ?

Grade:12

1 Answers

shabbir chatrissa
8 years ago
the initial momentum is zero.
let m1=3 kg m2=1 kg & their velocities be v1 & v2 respectively.
according to conservationh of momentum 
m1v+ m2v= 0
3v+ v= 0
v2 = -3v1
also,
 
0.5m1v22 + 0.5m2v2= 24000
m1v1+ m2v2= 4800
3v12 + v2= 4800
3v12 +(-3v1)= 4800
12v12 = 4800
v1 = 20
v2 = -60 (opp. direction)
KE of m2 is 1800.
 
I suggest you still cross check.